Series-parallel R,?24 3 2 R124 Ohm R2 = 30 Ohm R3 20 Ohm 3 2.6 6 4 5 In the circuit above a) Calculate the total resistance of the circuit
Answer #1

R2 and R3 are in parallel

Rp = [1/30 + 1/20]-1 = 12 ohms

Rp and R1 are in series

Requ = R1 + Rp

= 24 + 12

total resistance = 36 ohms